OpenStudy (nbarrera):

Help+ Medal! If x = 5 cm, find the area of this triangle A.675 cm2 B. 25 cm2 C.375 cm2 D.2700 cm2

Mehek (mehek14):

you need to ad an attachment

OpenStudy (nbarrera):

wait

OpenStudy (k_lynn):

The area formula of a triangle is \[\frac{ Base \times Height }{ 2 }\]Plug in the base and height and solve.

OpenStudy (k_lynn):

Also, do you have a picture you can attach so we can take a look at it?

OpenStudy (nbarrera):

ok there it is

Mehek (mehek14):

so plug in 5 6*5 and 9*5

OpenStudy (luffingsails):

1/2 base * height = 1/2 * 6*5 * 9*5 plug and chug.

OpenStudy (nbarrera):

so it's 30 and 45

OpenStudy (k_lynn):

Right, and that will give you the length of the base and height.

OpenStudy (k_lynn):

Once you do that, you would be able to solve.

Mehek (mehek14):

30*45*1/2

OpenStudy (nbarrera):

then i mutiply it

OpenStudy (nbarrera):

and =675?

OpenStudy (k_lynn):

Yep, you will multiply the base times the height.

OpenStudy (k_lynn):

Yep, @nbarrera 675 is your final answer. Great job.

OpenStudy (nbarrera):

thx

OpenStudy (k_lynn):

Sure, no problem.